# Scaling Laws for Task-Optimized Models of the Primate Visual Ventral Stream

### v1

@@ 1 Introduction
The advent of neural networks has revolutionized our understanding and modeling of complex neural processes. A particularly active area of study is the ventral visual stream in primates, a key pathway in the brain responsible for processing visual information Goodale & Milner (1992), Grill-Spector et al. (2001), Malach et al. (2002), Kriegeskorte et al. (2008). Neural networks, when trained on extensive datasets, have emerged as the most accurate quantitative tools for simulating the response patterns of neurons within this stream Yamins et al. (2014), Schrimpf et al. (2018). These advanced models offer a precise computational account of how neural mechanisms in the brain give rise to visual perception.
Recent developments in machine learning have emphasized the significance of both the volume of training data and the complexity of model architectures Kaplan et al. (2020), Hoffmann et al. (2022), Zhai et al. (2022), Bahri et al. (2022), Antonello et al. (2023), Muennighoff et al. (2023), Aghajanyan et al. (2023), Isik et al. (2024). These findings raise the question: Can we build better models of the brain by scaling up model architectures and dataset sizes? Recent studies have found that in pre-trained models, the number of parameters and dataset samples respectively seem to improve predictions of fMRI and behavioral measurements (Antonello et al., 2023, Muttenthaler et al., 2023). With the numerous differences between pre-trained models however, the relative contributions of model parameters and dataset size to brain and behavioral alignment are not clear.
In this paper, we examine how scaling – of model parameters and training dataset size – impacts the alignment of artificial neural networks with the primate ventral visual stream.
We systematically train models from a variety of architectural families on image classification datasets which allows us to independently control and observe the effects of model complexity and data volume. To capture the observed trends, we introduce parametric power-law trends that describe the impact of scale on alignment with behavior and brain regions along the visual ventral stream.
@@
Figure 1: a) For a given compute budget ($C$), we determine the scaling laws for maximal neural and behavioral alignment to the primate visual ventral stream. b) We find consistent scaling laws for brain and behavioral alignment across over 600 models. While we predict models to approach perfect behavioral alignment at large scales, the effect of scaling on brain alignment is already saturating.
@@ 2 Related Work
The ventral visual stream, a critical pathway in the primate brain, including humans, plays a key role in visual perception, extending from the occipital to the temporal lobes and serving as the ”what pathway” for object recognition and form representation Goodale & Milner (1992), Grill-Spector et al. (2001), Malach et al. (2002), Kriegeskorte et al. (2008). Beginning in the primary visual cortex (V1), where basic visual information from retinal ganglion cells is processed, the ventral stream proceeds through areas such as V2, V3, V4, and the inferotemporal cortex (IT), each responsible for increasingly complex features of visual perception Kandel et al. (2000). Despite decades of research and a wealth of brain data, the precise neural mechanisms underlying visual perception are not well understood.
@@
Particular artificial neural networks (ANNs) are the most accurate models of brain responses in the visual ventral stream and associated core object recognition behaviors Schrimpf et al. (2018; 2020). Models optimized for ecologically viable tasks (Yamins & DiCarlo, 2016) in particular have demonstrated strong brain and behavioral alignment (Yamins et al., 2014, Khaligh-Razavi & Kriegeskorte, 2014, Cadena et al., 2019, Schrimpf et al., 2018, Nayebi et al., 2018, Kietzmann et al., 2019, Rajalingham et al., 2018, Zhuang et al., 2021, Geiger et al., 2022) – notably these models are trained purely on image classification datasets, without fitting to brain data.
@@
Recent advancements in artificial intelligence are driven by scaling the model size and training data. Empirical evidence suggests a power-law relationship between model performance and both model parameters and dataset size, indicating that continued scaling will further improve performance Kaplan et al. (2020), Cherti et al. (2023), Zhai et al. (2022), Hoffmann et al. (2022), Dehghani et al. (2023), Henighan et al. (2020), Brown et al. (2020), Bahri et al. (2022), Hestness et al. (2017). The power-law exponents enable the optimal allocation of compute between model parameters and dataset samples, such that performance is maximized Kaplan et al. (2020), Hoffmann et al. (2022).
While scaling laws for machine learning performance has been extensively studied, the scaling laws for brain alignment remain unclear. Recent studies indeed suggest an involvement of both model size and data volume in the functional alignment with brain data Azabou et al. (2023), Benchetrit et al. (2023), Caro et al. (2024), Antonello et al. (2023). Conversely, Muttenthaler et al. Muttenthaler et al. (2023) indicate that sample size is critical for behavioral alignment. We here unify these results, in the realm of the primate visual ventral stream, into quantitative scaling laws for how model and dataset sizes relate to alignment with the brain and behavior.
@@ 3 Methods
We define the model’s alignment score $S$ (and an inverse Misalignment Score $L=1-S$) as the average across the V1, V2, V4, IT, and behavioral benchmark scores. Layers are committed to brain regions based on models trained on a full dataset, and applied to all variants trained with subsampled datasets. As we reused the same neural and behavioral data both to select the optimal model layer for readout and to assess the model’s alignment, we validated the benchmark results on a private split of each dataset on Brain-Score. We observed an almost perfect correlation between the results on the private and public splits (Appendix A).
@@
Following previous work on scaling laws Zhai et al. (2022), Hoffmann et al. (2022), Besiroglu et al. (2024), we fit power law functions in the form
@@
In that regard, following Hoffmann et al. (2022), Besiroglu et al. (2024) we fit a parametric function of the form
@@
where the loss ($\hat{L}$) is a function of parameter count ($N$) and number of samples seen ($D$). In Eq. 5, the first term represents the loss in an ideal data generation scenario (entropy), the second and the third terms reflect the under-performance of a model due to limitations in parameter and data size Hoffmann et al. (2022), Muennighoff et al. (2023). Following the example of Hoffmann et al. (2022), we learn variables $\{E,\;A,\;\alpha,\;B,\;\beta\}$ that characterizes misalignment by solving
@@
Using these learned variables, we predict the optimal allocation of compute ($C$) to $N$ and $D$ with a set equations:
@@
assuming that compute follows $C(N,D)\approx 6ND$ Kaplan et al. (2020), Hoffmann et al. (2022) relationship. However, we observe that this does not hold with different architectures, and various CNN families have a slightly different relationship of $C$, $N$, and $D$. As such, we assume a power-law relationship of the form
@@ 4.1 Scaling drives behavioral alignment, but saturates for neural alignment
Our experiments show a clear and consistent improvement in behavioral alignment as both model size and training dataset size increase. Fig 1.b illustrates this trend across different architectures and scaling axes. The curve $S=1-1.29\tilde{C}^{-0.07}$ converges to perfect alignment score of $1$ in the limit of $C$.
In contrast to behavioral alignment, neural alignment with specific brain regions demonstrated saturation as training compute scaled up in size. The curve represented by the formula $S=0.43-0.55\tilde{C}^{-0.07}$ represents a saturation at $0.43$. The diminishing returns in neural alignment imply that merely scaling up models and data is insufficient to achieve better alignment with higher-level neural representations.
@@ 4.3 More Data Is Better Than More Parameters
Figure 3: Scaling dataset size. a) Training on larger datasets improves brain alignment. Both ImageNet and Ecoset curves estimates ImageNet21k alignment quite closely, whereas datasets with lower-quality images (Places365 and WebVision) fall below the alignment scaling laws of higher-quality datasets. b) Model families with weaker inductive bias start at a lower alignment and require more data to improve.
@@
In contrast, expanding the model size without proportionally increasing the training data results in steeper diminishing returns in alignment performance. Larger models rapidly reach a point where additional parameters do not translate into meaningful improvements. Fig. 2c estimates a saturation level of $0.44$ by scaling model sizes with all samples of training data whereas Fig. 3a predicts maximum alignment of $0.48$ and $0.50$ for ImageNet and Ecoset respectively. This indicates that scaling training datasets overall improves brain alignment better than models scaling. Furthermore, Fig. 4b demonstrates that larger models of the same architecture family require much more samples to become s good.
Figure 4: Optimal compute allocation. a) Alignment as a function of both model and training dataset sizes. Marker size is log-proportional to model size. Compute should be spent XX/YY on model/dataset size respectively. b) Models start out at different alignments but converge to the same saturating point.
To quantitatively capture the joint interaction between data and model scaling, we fitted a parametric curve based on Eq.5, as shown in Fig.4a. This curve effectively models how compute ($C$), dataset size ($D$), and model size ($N$) collectively influence brain alignment. Utilizing the parametric relationships described in Eq. 9, we estimate that additional compute should be allocated following the scaling laws $D\approx C^{0.74}$ and $N\approx C^{0.26}$. These exponents indicate that, for optimal brain alignment, computational resources should be predominantly invested in increasing the dataset size rather than the model size.
@@ 5 Discussion
This disparity is further highlighted by the correlation between task performance and alignment depicted in Figure 6. Behavioral alignment closely tracks validation accuracy, improving hand-in-hand as models become more accurate. Consistent with prior work(Schrimpf et al., 2018, Linsley et al., 2023), neural alignment eventually saturates, indicating that factors other than task performance influence neural alignment.
Limitations.
@@
Second, we evaluated a subset of models focusing primarily on standard convolutional neural networks (e.g., ResNets, EfficientNets) and transformer-based architectures (e.g., ConvNeXts, ViTs). While these architectures cover a range of inductive biases and complexities, they do not encompass the full spectrum of possible neural network designs, such as recurrent networks or more biologically plausible models. We see scaling laws as an opportunity to extrapolate the alignment of models at scale, even if their current training is compute-constrained.
@@
Taken together, our results demonstrate that while scaling both model parameters and training data size enhances behavioral alignment with human visual perception, it leads to saturation in neural alignment with the primate ventral visual stream. Data scaling proves more effective than model scaling in improving alignment, emphasizing the critical role of extensive and diverse training datasets. We also find that architectural choices significantly influence alignment efficiency, with models possessing strong inductive biases—such as fully convolutional networks—achieving higher neural alignment even at initialization. Additionally, the impact of scaling varies across different brain regions, benefiting higher cortical areas more than early visual areas. These findings suggest that merely increasing scale is insufficient for modeling the intricate neural representations of the brain’s visual system. Future work should explore new approaches, including alternative architectures and training strategies, to develop models that more accurately reflect the complexities of neural processing in the primate visual cortex.
Figure 6: Correlation between task performance and alignment. a,b) Correlation between validation accuracy and brain (a) and behavioral (b) alignment. Behavioral alignment strongly correlates with task performance, whereas neural alignment shows a non-linear trend, reaching saturation. c) Pearson’s correlation coefficient per region, with all p-values less than $10^{-40}$
Appendix A Validation on private data
Figure 7: Private vs Public benchmarks

### v2

@@ 1 Introduction
The advent of neural networks has revolutionized our understanding and modeling of complex neural processes. A particularly active area of study is the ventral visual stream in primates, a key pathway in the brain responsible for processing visual information Goodale & Milner (1992); Grill-Spector et al. (2001); Malach et al. (2002); Kriegeskorte et al. (2008). Neural networks, when trained on extensive datasets, have emerged as the most accurate quantitative tools for simulating the response patterns of neurons within this stream Yamins et al. (2014); Schrimpf et al. (2018). These advanced models offer a precise computational account of how neural mechanisms in the brain give rise to visual perception.
Recent developments in machine learning have emphasized the significance of both the volume of training data and the complexity of model architectures Kaplan et al. (2020); Hoffmann et al. (2022); Zhai et al. (2022); Bahri et al. (2022); Antonello et al. (2023); Muennighoff et al. (2023); Aghajanyan et al. (2023); Isik et al. (2024). Henighan et al. (2020). These findings raise the question: Can we build better models of the brain by scaling up model architectures and dataset sizes? Recent studies have found that in pre-trained models, the number of parameters and dataset samples respectively seem to improve predictions of fMRI and behavioral measurements (Antonello et al., 2023; Muttenthaler et al., 2023). With the numerous differences between pre-trained models however, the relative contributions of model parameters and dataset size to brain and behavioral alignment are not clear.
In this paper, we examine how scaling – of model parameters and training dataset size – impacts the alignment of artificial neural networks with the primate ventral visual stream. We systematically train models from a variety of architectural families on image classification datasets which allows us to independently control and observe the effects of model complexity and data volume. To capture the observed trends, we introduce parametric power-law trends that describe the impact of scale on alignment with behavior and brain regions along the visual ventral stream.
@@
Figure 1: a) For a given compute budget ($C$), we determine the scaling laws for maximal neural and behavioral alignment to the primate visual ventral stream. b) We find consistent scaling laws for brain and behavioral alignment across over 600 models. While we predict models to approach perfect behavioral alignment of 1 at large scales, the effect of scaling on brain alignment is already saturating at 0.48.
@@ 2 Related Work
The ventral visual stream, a critical pathway in the primate brain, including humans, plays a key role in visual perception, extending from the occipital to the temporal lobes and serving as the ”what pathway” for object recognition and form representation Goodale & Milner (1992); Grill-Spector et al. (2001); Malach et al. (2002); Kriegeskorte et al. (2008). Beginning in the primary visual cortex (V1), where basic visual information from retinal ganglion cells is processed, the ventral stream proceeds through areas such as V2, V3, V4, and the inferotemporal cortex (IT), each responsible for increasingly complex features of visual perception Kandel et al. (2000).
Despite decades of research and a wealth of brain data, the precise neural mechanisms underlying visual perception are not well understood.
@@
Particular artificial neural networks (ANNs) are the most accurate models of brain responses in the visual ventral stream and associated core object recognition behaviors Schrimpf et al. (2018; 2020). Models optimized for ecologically viable tasks (Yamins & DiCarlo, 2016) in particular have demonstrated strong brain and behavioral alignment (Yamins et al., 2014; Khaligh-Razavi & Kriegeskorte, 2014; Cadena et al., 2019; Schrimpf et al., 2018; Nayebi et al., 2018; Kietzmann et al., 2019; Rajalingham et al., 2018; Zhuang et al., 2021; Geiger et al., 2022) – notably these models are trained purely on image classification datasets, without fitting to brain data.
@@
Recent advancements in artificial intelligence are driven by scaling the model size and training data. Empirical evidence suggests a power-law relationship between model performance and both model parameters and dataset size, indicating that continued scaling will further improve performance Kaplan et al. (2020); Cherti et al. (2023); Zhai et al. (2022); Hoffmann et al. (2022); Dehghani et al. (2023); Henighan et al. (2020); Brown et al. (2020); Bahri et al. (2022); Hestness et al. (2017). The power-law exponents enable the optimal allocation of compute between model parameters and dataset samples, such that performance is maximized Kaplan et al. (2020); Hoffmann et al. (2022).
While scaling laws for machine learning performance has been extensively studied, the scaling laws for brain alignment remain unclear. Recent studies indeed suggest an involvement of both model size and data volume in the functional alignment with brain data Azabou et al. (2023); Benchetrit et al. (2023); Caro et al. (2024); Antonello et al. (2023). Conversely, Muttenthaler et al. Muttenthaler et al. (2023) indicate that sample size is critical for behavioral alignment. We here unify these results, in the realm of the primate visual ventral stream, into quantitative scaling laws for how model and dataset sizes relate to alignment with the brain and behavior.
@@ 3 Methods
We define the model’s alignment score $S$ (and an inverse Misalignment Score $L=1-S$) as the average across the V1, V2, V4, IT, and behavioral benchmark scores. Layers are committed to brain regions based on models trained on a full dataset, and applied to all variants trained with subsampled datasets. As we reused the same neural and behavioral data both to select the optimal model layer for readout and to assess the model’s alignment, we validated the benchmark results on a private split of each dataset on Brain-Score. We observed an almost perfect correlation between the results on the private and public splits (Appendix C).
@@
Following previous work on scaling laws Zhai et al. (2022); Hoffmann et al. (2022); Besiroglu et al. (2024), we fit power law functions in the form
@@
In that regard, following Hoffmann et al. (2022); Besiroglu et al. (2024) we fit a parametric function of the form
@@
where the loss ($\hat{L}$) is a function of parameter count ($N$) and number of samples seen ($D$). In Eq. 5, the first term represents the loss in an ideal data generation scenario (entropy), the second and the third terms reflect the under-performance of a model due to limitations in parameter and data size Hoffmann et al. (2022); Muennighoff et al. (2023). Following the example of Hoffmann et al. (2022), we learn variables $\{E,\;A,\;\alpha,\;B,\;\beta\}$ that characterizes misalignment by solving
@@
Both Kaplan et al. (2020); Hoffmann et al. (2022) assume that compute follows the relationship $C(N,D)\approx 6ND$ to predict the optimal allocation of compute ($C$) to $N$ and $D$ using a set of equations with the learned variables mentioned above:
@@
However, we observe that $C(N,D)\approx 6ND$ does not hold with different architectures, and various CNN families have a slightly different relationship of $C$, $N$, and $D$. As such, we assume a power-law relationship of the form
@@ 4.1 Scaling drives behavioral alignment, but saturates for neural alignment
Our experiments show a clear and consistent improvement in behavioral alignment as both model size and training dataset size increase. Fig 1.b illustrates this trend across different architectures and scaling axes. The curve $S=1-1.4\tilde{C}^{-0.06}$ converges to perfect alignment score of $1$ in the limit of $C$.
In contrast to behavioral alignment, neural alignment with specific brain regions demonstrated saturation as training compute scaled up in size. The curve represented by the formula $S=0.48-0.55\tilde{C}^{-0.16}$ represents a saturation at $0.48$. The diminishing returns in neural alignment imply that merely scaling up models and data is insufficient to achieve better alignment with higher-level neural representations.
@@ 4.3 More Data Is Better Than More Parameters
Figure 3: Scaling Dataset Size. a) Training on larger datasets enhances brain alignment. The alignment scaling curves derived from ImageNet and EcoSet closely estimate the alignment achieved when using ImageNet21k. In contrast, datasets with specialized image distributions—such as Places365—fall below the alignment scaling laws established by these generalist datasets. b) Model families with weaker inductive bias start at a lower alignment and require more data to improve.
@@
In contrast, expanding the model size without proportionally increasing the training data results in steeper diminishing returns in alignment performance. Larger models rapidly reach a point where additional parameters do not translate into meaningful improvements. Fig. 2c estimates a saturation level of $0.44$ by scaling model sizes with all samples of training data whereas Fig. 3a predicts maximum alignment of $0.48$ and $0.50$ for ImageNet and Ecoset respectively. This indicates that scaling training datasets overall improves brain alignment better than models scaling. Furthermore, Fig. 4b demonstrates that larger models of the same architecture family require much more samples to achieve the same level of alignment.
Figure 4: Optimal Compute Allocation. a) Alignment as a function of both model and training dataset sizes. Marker size is log-proportional to model size. Compute should be spent 0.3/0.7 on model/dataset size respectively. b) Models start out at different alignments but converge to the same saturating point.
To quantitatively capture the joint interaction between data and model scaling, we fitted a parametric curve based on Eq.5, as shown in Fig.4a. This curve effectively models how compute ($C$), dataset size ($D$), and model size ($N$) collectively influence brain alignment. Utilizing the parametric relationships described in Eq. 9, we estimate that additional compute should be allocated following the scaling laws $D\approx C^{0.7}$ and $N\approx C^{0.3}$. These exponents indicate that, for optimal brain alignment, computational resources should be predominantly invested in increasing the dataset size rather than the model size.
@@ 5 Discussion
This disparity is further highlighted by the correlation between task performance and alignment depicted in Figure 6. Behavioral alignment closely tracks validation accuracy, improving hand-in-hand as models become more accurate. Consistent with prior work(Schrimpf et al., 2018; Linsley et al., 2023), neural alignment eventually saturates, indicating that factors other than task performance influence neural alignment.
Generalization Beyond Supervised Training.
We assessed whether alternative training paradigms can overcome the limitations observed in neural alignment under supervised learning. Figure 7a illustrates the scaling of alignment as a function of compute spent during self-supervised training of ResNet models using SimCLR Chen et al. (2020) on ImageNet. The results confirm the trends observed in supervised training: behavioral alignment continues to improve with increased compute, following a strong power-law relationship, while neural alignment approaches a saturation point. This consistency suggests that the saturation in neural alignment is not exclusive to supervised learning but may be inherent to the models or datasets employed.
The region-specific breakdown (as illustrated in Figure 13 of the appendix) further reinforces this observation. Even in a self-supervised learning context, higher-level visual areas like IT and behavioral outputs demonstrate more pronounced improvements with increased compute, while early visual areas like V1 and V2 show minimal gains. This suggests that the hierarchical nature of neural alignment is a fundamental characteristic that transcends specific training methodologies.
Additionally, we explored the impact of adversarial fine-tuning on alignment performance. In Figure 7b, ResNet models trained on subsets of ImageNet were fine-tuned adversarially for 10 epochs using the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015; Wong et al., 2020). Importantly, the scaling curves were estimated solely from the non-adversarial runs, yet the adversarially fine-tuned models exhibited improvements along these existing scaling curves. This indicates that adversarial training can enhance alignment without deviating from the established scaling behavior.
Figure 6: Correlation between Task Performance and Alignment. a,b) Correlation between validation accuracy (ImageNet & EcoSet) and brain (a) and behavioral (b) alignment. Behavioral alignment strongly correlates with task performance, whereas neural alignment shows a non-linear trend, reaching saturation. c) Pearson’s correlation coefficient per region, with all p-values less than $10^{-40}$.
Impact of Architectural Inductive Biases on Alignment Dynamics.
Our evaluation of alignment during training reveals that the alignment behavior varies significantly across different model architectures. Figure 7.c shows that while various models eventually converge to similar alignment levels with sufficient training, fully convolutional architectures—such as ResNets and EfficientNets—exhibit substantially higher alignment scores at the very beginning of training. This early advantage suggests that these architectures possess inherent features that align closely with neural data from the primate ventral visual stream even before learning from data occurs.
Further analysis in Figure 10 of the appendix confirms that this initial high alignment is due to the strong inductive biases present in fully convolutional networks. These biases enable the models to start with representations already well-suited for neural alignment. Figure 11 in the appendix reinforces this finding by demonstrating that models with strong inductive biases achieve higher initial alignment compared to architectures like ConvNeXt and ViT, which have weaker inductive biases.
Influence of Learning Signals on Alignment Dynamics.
Our investigation reveals that the type of learning signal plays a crucial role in the dynamics of alignment during training. Figure 7d illustrates the alignment trajectories of ResNet50 and ViT-S models trained on ImageNet using supervised learning, SimCLR, and DINO (Caron et al., 2021) methods. Notably, the ViT-S model requires significantly more training steps to achieve the same level of alignment under supervised learning compared to when trained with self-supervised objectives like DINO and SimCLR. In contrast, the ResNet50 model, which possesses strong inductive biases due to its convolutional architecture, exhibits relatively consistent alignment dynamics across different learning signals. This robustness implies that models with strong inductive biases are less affected by the choice of training objective, whereas architectures like ViT-S benefit more substantially from rich, self-supervised feedback to achieve optimal alignment.
Limitations and Future Directions.
@@
Second, we evaluated a subset of models focusing primarily on standard and modern convolutional neural networks (e.g., ResNets and ConvNeXts), transformer-based architectures (e.g. ViTs) and recurrent networks (CORnet-S). While these architectures cover a range of inductive biases and complexities, they do not encompass the full spectrum of possible neural network designs, such as more biologically plausible models. We see scaling laws as an opportunity to extrapolate the alignment of models at scale, even if their current training is compute-constrained.
@@
Taken together, our results demonstrate that while scaling both model parameters and training data size enhances behavioral alignment with human visual perception, it leads to saturation in neural alignment with the primate ventral visual stream. Data scaling proves more effective than model scaling in improving alignment, emphasizing the critical role of extensive and diverse training datasets. We also find that architectural choices significantly influence alignment efficiency, with models possessing strong inductive biases—such as fully convolutional networks—achieving higher neural alignment even at initialization. Additionally, the impact of scaling varies across different brain regions, benefiting higher cortical areas more than early visual areas. These findings suggest that merely increasing scale is insufficient for modeling the intricate neural representations of the brain’s visual system. Future work should investigate new approaches, including alternative architectures and training strategies, to develop models that more accurately reflect the complexities of neural processing in the primate visual cortex.
To push neural alignment beyond current saturation levels, future research should explore adversarial training methods that encourage models to learn more robust, brain-like representations. Leveraging biologically inspired architectures such as VOneNets (Dapello et al., 2020) may lead to more compute-efficient models achieving higher neural alignment without extensive scaling. Additionally, investigating co-training with brain data—integrating neural recordings directly into the training process—could enhance both neural and behavioral alignment, paving the way for more accurate and efficient brain-like models.
Figure 7: Alignment Scaling of Alternative Training Strategies. a) Unsupervised: ResNet models trained with SimCLR improve behavioral alignment with compute, while neural alignment saturates. b) Adversarial Robustness: Fine-tuning ResNets with adversarial training (FGSM) enhances alignment along the scaling curve. c) Architectural Prior: ResNets and EfficientNets exhibit higher initial alignment in early phases of training, due to strong inductive biases, unlike ConvNeXt and ViT. d) Alignment dynamics vary with training objectives, but converge to the same alignment saturation.
Appendix A Implementation Details
Our experiments are conducted using the PyTorch framework (Paszke et al., 2019), with Composer (Team, 2021) employed as the GPU orchestration tool to efficiently manage computational resources.
For image augmentations, we leverage the Albumentations Buslaev et al., 2020 library due to its rich set of augmentation techniques, which are crucial for enhancing model robustness and preventing overfitting. In experiments involving self-supervised learning, we use the Lightly Susmelj et al. (2020) library to facilitate the implementation of self-supervised losses, augmentations, and model heads. This library streamlines the process of setting up models for SimCLR and DINO training methods.
To generate adversarial examples for adversarial fine-tuning, we employ the Torchattacks library Kim (2020). Specifically, we use the Fast Gradient Sign Method (FGSM) to create perturbations that challenge the models, aiming to enhance their alignment with neural representations by exposing them to adversarial inputs.
Appendix B Additional Image Datasets
To further validate our findings across diverse image distributions and to estimate scaling curves across different sample scales, we trained ResNet18 models on subsets of several large-scale image datasets: ImageNet-21k-P, WebVision-P, iNaturalist, and Places365. Below, we provide detailed descriptions of each dataset.
B.1 ImageNet21k-P
ImageNet-21k-P is a processed subset of the full ImageNet-21k dataset (Ridnik et al., 2021), which originally contains over 14 million images organized into more than 21,000 categories following the WordNet hierarchy. The ”P” denotes a pruned version where classes with insufficient images or noisy labels are filtered out to enhance dataset quality. This results in a refined dataset that maintains the richness of the original ImageNet-21k while improving label accuracy and image relevance. The resulting dataset contains approximately 11 million training images across 10,450 classes.
B.2 WebVision-P
The WebVision dataset (Li et al., 2017) is a large-scale web image dataset designed to provide a real-world, noisy alternative to ImageNet. It originally contains over 16 million images categorized into 5,000 classes. The images are collected from the internet using queries from search engines like Google and Flickr, leading to a dataset that includes label noise, varying image resolutions, and diverse visual contexts. Due to classes with very few available samples, we processed the WebVision dataset similarly to ImageNet-21k-P to remove classes with insufficient images. The resulting dataset, which we denote as WebVision-P, contains approximately 13.5 million training images across 4,189 categories.
B.3 iNaturalist
iNaturalist Van Horn et al. (2018) contains 2.7 million photographs of organisms in their natural environments, representing 10000 species. The dataset features highly specialized fine-grained categories and natural backgrounds, offering insight into how domain-specific visual features influence alignment scaling.
B.4 Places365
Places365 Zhou et al. (2017) is a large-scale scene-centric dataset containing approximately 1.8 million training images across 365 scene categories. Unlike object-centric datasets such as ImageNet, Places365 focuses on the recognition of environmental scenes, including natural landscapes, urban settings, and indoor environments. Each category includes a wide variety of images to capture the diversity within scene types.
B.5 infiMNIST
The MNIST dataset (LeCun et al., 1998) is a classic benchmark in machine learning, comprising 70,000 grayscale images of handwritten digits (0-9), each sized 28×28 pixels. To expand this dataset for more extensive experimentation, we utilize the Infinite MNIST (Infimnist) tool Loosli et al., 2007, which generates additional MNIST-like samples through data augmentation techniques. We create an extended dataset by modifying the original training dataset 19 additional times, resulting in a total of 1.2 million images. This enlarged dataset allows for a more thorough evaluation of scaling effects on the alignment.
Appendix C Validation on private data
Figure 8: Public benchmarks used in this study correlates highly with private benchmarks on Brain-Score